Please help me solve the following equations
X+3y=4
2x-y=8

Answer by longjonsilver(2297)

Look at my Lessons on Simultaneous equation, on this website... they help.

solution is:
x+3y=4 --eqn1
2x-y=8 --eqn2

scale up eqn1 by factor 2, giving:
2x+6y=8 --eqn1
2x-y=8 --eqn2

Now, the x-terms are the same digit...both a 2. Looking at the signs of these, they are the SSSSSSSSSame (both +), so we SSSSSSSSubtract the equations, to give:

5y = 0
so y = 0.

Hence, x+3y=4 becomes
x+3(0) = 4
x = 4

and CHECK in the other equation:
2x-y=8 becomes
2(4)-0 = 8
8=8 --> which is correct

jon.